Find the EAR in each of the following cases. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16. Use 365 days in a year.) 


Stated Rate (APR)Number of Times CompoundedEffective Rate (EAR)


10.50%Quarterly
%
15.50%Monthly
%
16.50%Daily
%
12.50%Semiannually
%
Answer #1
EAR = ( 1 + r )^n - 1
Compounded Quarterly
EAR= ( 1 + 10.50%/4)^4 - 1
EAR= 10.92%
Compounded Monthly
EAR= ( 1 + 15.50%/12)^12 - 1
EAR= 16.65%
Compounded Daily
EAR= ( 1 + 16.50%/365)^365 - 1
EAR= 17.93%
Compounded Semianually
EAR= ( 1 + 12.50%/2)^2 - 1
EAR= 12.89%